# Leap: molecular synthesisability scoring with intermediates

### v1

@@ 1 Introduction
The ability to predict whether a compound can be made is key to accelerating drug discovery. Generative methods that neglect synthesisability often suggest compounds that are challenging or impossible to make Gao and Coley (2020). Attempting to synthesise these problematic compounds is costly and time-consuming. Therefore, it is crucial to assess synthesisability as part of the scoring functions applied within generative methods to prioritise synthetically accessible compounds.
Furthermore, the concept of synthesisability should be regarded as dynamic and dependent on the compounds available at a specific time. Synthesis planning algorithms Gao et al. (2018); Genheden et al. (2020); ibm ; Segler et al. (2018) are used to identify possible synthetic pathways for specific molecules by using readily accessible compounds. However, these methods are computationally expensive and slow to apply exhaustively within generative workflows, which can involve the assessment of hundreds of thousands to millions of compounds Gao and Coley (2020). Therefore, it makes them unsuitable to assess synthesisability when speed is required to provide feedback to generative models. In practice, this limits the applicability of generative workflows to real world scenarios where an fast, automated and accurate metric of chemical tractability is critical to enable chemists to prioritise high quality compounds that minimise development time and cost.
To address this challenge, scoring methods such as SAScore Ertl and Schuffenhauer (2009), SCScore Coley et al. (2018), and RAScore Thakkar et al. (2021) were developed to assess the synthetic accessibility of molecules. SAScore assumes that the complexity of synthesising a molecule is correlated with the scarcity of its fragments in databases of bioactive compounds (e.g. PubChem Kim et al. (2022) or ChEMBL Gaulton et al. (2011)) and checks for complex chemistry like stereocenters and ring systems. SCScore is designed to ensure that, on average, the products of chemical reactions exhibit greater synthetic complexity compared to their reactants. RAScore Thakkar et al. (2021) classifies whether a synthetic route can be identified for a particular compound or not by the retrosynthesis tool AiZynthFinder Genheden et al. (2020). During inference, none of these methods are aware of possible available intermediates for the query molecules. In drug discovery projects, a common strategy is to identify, through the synthesis of molecules around a specific chemical space, large quantities of useful synthesisable intermediates from which further compounds can be readily made Hughes et al. (2011). When failing to adapt to this information, existing scorers can overestimate the synthetic complexity of molecules produced in this way Makara et al. (2021), hindering their ability to identify compounds as synthesisable, particularly when these intermediates are complex.
@@ 2.1 Problem Setting
Our method is based on the success of transformers Vaswani et al. (2023) in predicting single-step forward reactions (reactants to product), as well as single-step retrosynthesis (product to reactants)  Schwaller et al. (2019); Wang et al. (2021) by extending the approach to multi-step routes. We achieve this by encoding synthesis routes as strings, and training a transformer model (GPT-2 Radford et al. (2019)) to predict them. To guide the model towards the expected route, we prompt it with the expected building blocks. Finally, we fine-tune this model using a regression head to predict tree depth for target molecules.
@@ 2.2.1 Data Representation
Each molecule is encoded as Smiles. Inspired by the reaction smiles smi , we represent these in a product>>reactants format, where each reactant is separated by a dot. Reactions occurring at the same depth are separated by a ; , while we use : to separate reactions at different depths of the tree.
@@ 2.2.2 Model Architecture and Training
Our model is a GPT-2 model with 8 attention heads and 5 hidden layers. We use HuggingFace’s implementation Wolf et al. (2020). The hidden states have a dimension of 512, while the feed-forward layers have a dimension of 1024. We train the model for 16 epochs, use AdamW optimiser with a learning rate of 0.0003, and a cosine learning rate scheduler with 2000 warmup steps. We train the model on a language modelling task, where the goal is to predict the next token.
@@ 2.2.3 Dataset
To pre-train our model, we use a combination of two datasets. As no dataset of multi-step synthetic routes exists, methods have been developed to generate pathways from existing single-step reaction datasets. The first dataset from  Liu et al. (2023), uses dynamic programming to extract synthesis routes from patent reactions found within the USPTO-all Lowe (2012) dataset. We use train, test and validation splits from Liu et al. (2023) which gives 39717, 5479 and 5493 routes for each.
The dataset from Liu et al. (2023) is skewed both towards short routes of an average depth of three, and routes that only extend down a single branch of the synthesis tree (i.e. linear rather than convergent synthesis routes). To increase diversity, and allow for longer and more branched routes, we extend the training set with a second dataset based on Gao et al. (2022). This models the problem of synthetic tree generation as a Markov decision process with a random policy over reactants and reaction templates. Reaction templates are rule-based patterns extracted from existing reactions to define transformations on molecules that represent valid chemical reactions. From this we obtain 116029 synthesis routes, which we add to the training set.
We convert all routes to the string format described in Section 2.2.1. We augment the data by randomising the Smiles as suggested in Arús-Pous et al. (2019), and randomly permuting both the reactions located at the same depth of the route, and the building blocks in the prompt. We apply augmentation three times per route.
@@ 2.3.3 Dataset
The fine-tuning dataset is a sample of 290418 ChEMBL Gaulton et al. (2011) molecules that we use to train our model. In addition, we curate a held-out dataset of 4989 project molecules coming directly from internal drug discovery projects, to challenge the model on out-of-domain prediction.
We use AiZynthFinder Genheden et al. (2020) with default parameters to compute the routes of both datasets. A route is considered to be solved if all leaves are purchasable molecules. If more than one route is found for a molecule, the route scored the highest by AiZynthFinder is retained. For solved routes, we compute the depth of the tree. Since the maximum tree depth allowed by AiZynthFinder’s default parameters is 7, we assign a depth of 10 for molecules where a route is not found.
@@ 3.1 Identifying synthesisable molecules
To determine molecular synthesisability we compute routes using AiZynthFinder Genheden et al. (2020) for both test and project molecules. If a route can be found, the molecule is classified as synthesisable.
To permit comparison between Leap and RAScore, SAScore and SCScore, and to establish a baseline for further experiments, we do not consider intermediates in this experiment. We score all molecules with all scorers, and use these to compute the ROC-AUC curves shown in Figure 2. Leap outperforms all other methods in terms of AUC score for both test and project molecules. This demonstrates that our model remains useful even when intermediates are not present. Leap also shows superior performance on the hold-out set of project molecules, showing its ability to generalise beyond its training domain. Conversely, even though RAScore was also trained on ChEMBL molecules, its predictive performance deteriorates on out-of-domain molecules as previously observed in Thakkar et al. (2021).
@@ 3.4 Comparison to baseline models
To validate our architecture and our pre-training method, we conduct experiments using three baseline models: CatBoost, GPT-2 (i.e. Leap with no pre-training) and BERT Devlin et al. (2018). We compare their performance on R2, mean absolute error (MAE) and mean squared error (MSE). For CatBoost, intermediate and target molecules are converted to ECFP4 fingerprints, and then concatenated. For BERT, they are passed in as two sentences, separated by SEP token. The results are reported in Table 1. We show that pre-training helped boost the performance of Leap, and that the fine-tuned Leap outperforms both CatBoost and BERT on all the metrics.

### v2

@@ 1 Introduction
The ability to predict whether a compound can be made is key to accelerating drug discovery. Generative methods that neglect synthesisability often suggest compounds that are challenging or impossible to make [8]. Attempting to synthesise these problematic compounds is costly and time-consuming. Therefore, it is crucial to assess synthesisability as part of the scoring functions applied within generative methods to prioritise synthetically accessible compounds.
Furthermore, the concept of synthesisability should be regarded as dynamic and dependent on the compounds available at a specific time. Synthesis planning algorithms [7, 11, 1, 19] are used to identify possible synthetic pathways for specific molecules by using readily accessible compounds. However, these methods are computationally expensive and slow to apply exhaustively within generative workflows, which can involve the assessment of hundreds of thousands to millions of compounds [8]. Therefore, it makes them unsuitable to assess synthesisability when speed is required to provide feedback to generative models. In practice, this limits the applicability of generative workflows to real world scenarios where an fast, automated and accurate metric of chemical tractability is critical to enable chemists to prioritise high quality compounds that minimise development time and cost.
To address this challenge, scoring methods such as SAScore [6], SCScore [4], and RAScore [20] were developed to assess the synthetic accessibility of molecules. SAScore assumes that the complexity of synthesising a molecule is correlated with the scarcity of its fragments in databases of bioactive compounds (e.g. PubChem [13] or ChEMBL [10]) and checks for complex chemistry like stereocenters and ring systems. SCScore is designed to ensure that, on average, the products of chemical reactions exhibit greater synthetic complexity compared to their reactants. RAScore [20] classifies whether a synthetic route can be identified for a particular compound or not by the retrosynthesis tool AiZynthFinder [11]. During inference, none of these methods are aware of possible available intermediates for the query molecules. In drug discovery projects, a common strategy is to identify, through the synthesis of molecules around a specific chemical space, large quantities of useful synthesisable intermediates from which further compounds can be readily made [12]. When failing to adapt to this information, existing scorers can overestimate the synthetic complexity of molecules produced in this way [16], hindering their ability to identify compounds as synthesisable, particularly when these intermediates are complex.
@@ 2.1 Problem Setting
Our method is based on the success of transformers [21] in predicting single-step forward reactions (reactants to product), as well as single-step retrosynthesis (product to reactants)  [18, 22] by extending the approach to multi-step routes. We achieve this by encoding synthesis routes as strings, and training a transformer model (GPT-2 [17]) to predict them. To guide the model towards the expected route, we prompt it with the expected building blocks. Finally, we fine-tune this model using a regression head to predict tree depth for target molecules.
@@ 2.2.1 Data Representation
Each molecule is encoded as Smiles. Inspired by the reaction smiles [2], we represent these in a product>>reactants format, where each reactant is separated by a dot. Reactions occurring at the same depth are separated by a ; , while we use : to separate reactions at different depths of the tree.
@@ 2.2.2 Model Architecture and Training
Our model is a GPT-2 model with 8 attention heads and 5 hidden layers. We use HuggingFace’s implementation [23]. The hidden states have a dimension of 512, while the feed-forward layers have a dimension of 1024. We train the model for 16 epochs, use AdamW optimiser with a learning rate of 0.0003, and a cosine learning rate scheduler with 2000 warmup steps. We train the model on a language modelling task, where the goal is to predict the next token.
@@ 2.2.3 Dataset
To pre-train our model, we use a combination of two datasets. As no dataset of multi-step synthetic routes exists, methods have been developed to generate pathways from existing single-step reaction datasets. The first dataset from  Liu et al. [14], uses dynamic programming to extract synthesis routes from patent reactions found within the USPTO-all [15] dataset. We use train, test and validation splits from Liu et al. [14] which gives 39717, 5479 and 5493 routes for each.
The dataset from Liu et al. [14] is skewed both towards short routes of an average depth of three, and routes that only extend down a single branch of the synthesis tree (i.e. linear rather than convergent synthesis routes). To increase diversity, and allow for longer and more branched routes, we extend the training set with a second dataset based on Gao et al. [9]. This models the problem of synthetic tree generation as a Markov decision process with a random policy over reactants and reaction templates. Reaction templates are rule-based patterns extracted from existing reactions to define transformations on molecules that represent valid chemical reactions. From this we obtain 116029 synthesis routes, which we add to the training set.
We convert all routes to the string format described in Section 2.2.1. We augment the data by randomising the Smiles as suggested in Arús-Pous et al. [3], and randomly permuting both the reactions located at the same depth of the route, and the building blocks in the prompt. We apply augmentation three times per route.
@@ 2.3.3 Dataset
The fine-tuning dataset is a sample of 290418 ChEMBL [10] molecules that we use to train our model. In addition, we curate a held-out dataset of 4989 project molecules coming directly from internal drug discovery projects, to challenge the model on out-of-domain prediction.
We use AiZynthFinder [11] with default parameters to compute the routes of both datasets. A route is considered to be solved if all leaves are purchasable molecules. If more than one route is found for a molecule, the route scored the highest by AiZynthFinder is retained. For solved routes, we compute the depth of the tree. Since the maximum tree depth allowed by AiZynthFinder’s default parameters is 7, we assign a depth of 10 for molecules where a route is not found.
@@ 3.1 Identifying synthesisable molecules
To determine molecular synthesisability we compute routes using AiZynthFinder [11] for both test and project molecules. If a route can be found, the molecule is classified as synthesisable.
To permit comparison between Leap and RAScore, SAScore and SCScore, and to establish a baseline for further experiments, we do not consider intermediates in this experiment. We score all molecules with all scorers, and use these to compute the ROC-AUC curves shown in Figure 2. Leap outperforms all other methods in terms of AUC score for both test and project molecules. This demonstrates that our model remains useful even when intermediates are not present. Leap also shows superior performance on the hold-out set of project molecules, showing its ability to generalise beyond its training domain. Conversely, even though RAScore was also trained on ChEMBL molecules, its predictive performance deteriorates on out-of-domain molecules as previously observed in Thakkar et al. [20].
@@ 3.4 Comparison to baseline models
To validate our architecture and our pre-training method, we conduct experiments using three baseline models: CatBoost, GPT-2 (i.e. Leap with no pre-training) and BERT [5]. We compare their performance on R2, mean absolute error (MAE) and mean squared error (MSE). For CatBoost, intermediate and target molecules are converted to ECFP4 fingerprints, and then concatenated. For BERT, they are passed in as two sentences, separated by SEP token. The results are reported in Table 1. We show that pre-training helped boost the performance of Leap, and that the fine-tuned Leap outperforms both CatBoost and BERT on all the metrics.